OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 4 and 9. What is the length of the altitude? 9 13 4 6

OpenStudy (anonymous):

@Rav

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

do u know this? @terenzreignz

terenzreignz (terenzreignz):

Yeah. But I think I'll wait for @Rav to have his/her say :)

OpenStudy (anonymous):

which one do u think it is? i think its 13 or 4

terenzreignz (terenzreignz):

You can't have two answers.

OpenStudy (anonymous):

im preety sure its 13 though,is it?

terenzreignz (terenzreignz):

No.

OpenStudy (anonymous):

ughhh so its 4?

terenzreignz (terenzreignz):

If I suspect you're guessing, I'll just keep saying no even if you mention the correct answer :P Best not to guess :)

OpenStudy (anonymous):

i just dont really know the answer,can u give it to me and explain how u got it?

terenzreignz (terenzreignz):

What does @Rav think? :)

OpenStudy (anonymous):

i dont think his gonna come. i wanna get through this quick since its late and learn it

terenzreignz (terenzreignz):

This is not an exam, right?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

So... what do you know about altitudes of right triangles?

OpenStudy (anonymous):

idk..

terenzreignz (terenzreignz):

Well, we have to work with what we know... Altitudes drawn to hypotenuses make the two resulting triangles similar to each other.

OpenStudy (anonymous):

yeahh!

terenzreignz (terenzreignz):

Well, what does that mean, when two triangles are similar?

OpenStudy (anonymous):

when they are alike

terenzreignz (terenzreignz):

alike... but specifically?

OpenStudy (anonymous):

when they are congruent or have something identical

terenzreignz (terenzreignz):

Congruent triangles are certainly similar. Can you expound on "have something identical"?

OpenStudy (anonymous):

like the same lines.

OpenStudy (anonymous):

can u just help me this one? The picture shows a portion of an irrigation system for a farm. QR is the main water pipe. http://oi49.tinypic.com/mwvcbm.jpg What is the length of the water pipe PR? 10 feet 20 feet 14.1 feet 28.3 feet I dont get it @terenzreignz

OpenStudy (anonymous):

Start with what you know. You know SOHCAHTOA: sin = O/H, cos = A/H, tan = O/A You know one angle, 45 degrees You know one side, 20 ft, adjacent to the angle. What do you need, sin cos or tan?

OpenStudy (anonymous):

im not sure @Rav

OpenStudy (anonymous):

Look at what you have then. I can't give you the answer here. I've walked you through a few trigonometry questions and I want to make sure you're understanding the material. Math problems are ALL about selecting the right equation or correlation based on what information you already have.

OpenStudy (anonymous):

ik and i appreciate that. Can you give me the answer and explain how you got it then? i understand really good that way. cuz idk anything about this

OpenStudy (anonymous):

Well in this problem you know the longest side (20 ft) and you want the length of the side next to the angle (the adjacent side). So what you want is a relationship that combines the hypotenuse (H) and the adjacent (A) into one formula. You can choose sin (O/H), cos (A/H) or tan (O/A). Which one should you use?

OpenStudy (anonymous):

umm tan?

OpenStudy (anonymous):

Plug in the values. You want to find A and you know H. So using tan: tan 45 = O/20 20 tan 45 = O Does that help?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

guide me to the answer so i know how to work out everything! it helps alot!

OpenStudy (anonymous):

Remember your options. tan 45 = O/A A tan 45 = O sin 45 = O/20 20 sin 45 = O cos 45 = O/20 20 cos 45 = A Which equation gives you the answer you want? And by that I mean, which one has only a number on one side and the unknown variable (A) on the other?

OpenStudy (anonymous):

That last one should read cos 45 = A/20 20 cos 45 = A

OpenStudy (anonymous):

Cosec 45

OpenStudy (anonymous):

Cosec 45 = 1/sin 45 = H/O = 20/O Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If that's correct, then tell me the value of A.

OpenStudy (anonymous):

Oh not its not correct. but idk which one is correct.

OpenStudy (anonymous):

Why don't you know? What do we need to work on first?

OpenStudy (anonymous):

im just verry tirred and for now i just want to know the answer and then tomorrow actually try to learn it. are u going to get on tomorrow?

OpenStudy (anonymous):

Yes I'll be on here tomorrow. The answer is 10.5 ft. cos 45 = A/H = A/20 20 cos 45 = A A = 10.5

OpenStudy (anonymous):

theres no 10.5.. only 10. @rav

OpenStudy (anonymous):

Then it must be 10, but remember the two sides always add up to more than the hypotenuse, so it can't be exactly 10.

OpenStudy (anonymous):

Ok, can u give me the answer to this one too? Tomorrow imma get on so i can learn all this. Two end points of a line segment are (-12, -2) and (-6, -10). What are the coordinates of the point on the line through which its bisector passes? (-18, -12) (-9, -6) (-7, -8) (-11, -4)

OpenStudy (dan815):